OpenStudy (rayman5766):

which is a solution's to the equation 5x + 2y =9 last one easy

OpenStudy (rayman5766):

help me please

OpenStudy (daniellelovee):

1 Subtract 2y from both sides 5x=9−2y 5x=9−2y 2 Divide both sides by 5 x=9−2y/5

OpenStudy (daniellelovee):

whattt xD

OpenStudy (welshfella):

its best to enclose the 9 - 2y in parentheses to avoid confusion y = (9 - 2y) / 5

OpenStudy (daniellelovee):

from there you can fin the y and the x

OpenStudy (welshfella):

* x = (9 - 2y) / 5

OpenStudy (rayman5766):

oh is see that makes it easier

OpenStudy (daniellelovee):

well give us your options please and thank you for pointing that out :)

OpenStudy (rayman5766):

wait so is it

OpenStudy (rayman5766):

x = -2 y + 9/5

OpenStudy (daniellelovee):

yes that can be

OpenStudy (rayman5766):

no that cant' be it because that's not one of my options

OpenStudy (daniellelovee):

then show me your options please

OpenStudy (rayman5766):

A (3,3) B (1,2) C (0,5) D (-2,2)

OpenStudy (daniellelovee):

oh then just replace the numbers in the equation and find one that equals 9

OpenStudy (welshfella):

again its best not to write it that way because the whole part, -2y + 9, is divided by 5. The way its written it can be interpreted that only the 9 is divided by 5 . That's why you should write it as (-2y + 9) / 5.

OpenStudy (rayman5766):

A

OpenStudy (rayman5766):

3*2 is 9

OpenStudy (rayman5766):

3*3 is 9 not 2

OpenStudy (daniellelovee):

dont guess show your works example A (3,3) 5(3)+2(3)=21 21 is not 9

OpenStudy (rayman5766):

right i knew that :)

OpenStudy (daniellelovee):

lol alright then show your work :)

OpenStudy (rayman5766):

5(1)+2(2)=14

OpenStudy (daniellelovee):

and 14 is not 9 so no carry on :)

OpenStudy (rayman5766):

is it D

OpenStudy (rayman5766):

because C is 10

OpenStudy (rayman5766):

you there

OpenStudy (daniellelovee):

sorry i left im back

OpenStudy (daniellelovee):

and its B because 5(2)+2(2)=9

OpenStudy (rayman5766):

darn i had that before

OpenStudy (daniellelovee):

lol

OpenStudy (rayman5766):

wait thats D

OpenStudy (daniellelovee):

I just said what it was

OpenStudy (rayman5766):

you said B wic is 1,2

OpenStudy (daniellelovee):

correct

OpenStudy (rayman5766):

si is it b or DDDDDD

OpenStudy (daniellelovee):

i already said

OpenStudy (daniellelovee):

A (3,3) B (1,2) C (0,5) D (-2,2)

OpenStudy (rayman5766):

which one D has 2,2 B has 1,2 so reap this to me

OpenStudy (daniellelovee):

read what I wrote please I gave you the answer

OpenStudy (rayman5766):

so its D

OpenStudy (rayman5766):

can you reapt the answer

OpenStudy (daniellelovee):

you tell me I gave you my work already I don't know you guess because i told you the answer

OpenStudy (rayman5766):

all you have to do is reapt the answer